# Thermodynamic Gravity with Non-Extensive Horizon Entropy and Topological Calibration

### v1

@@ 1 Introduction
A recent framework proposed by Lü, Di Gennaro and Ong  et al. (2025) takes an important step in this direction. In particular, working in a local-horizon framework in the spirit of Jacobson, they replace the Bekenstein–Hawking entropy by a Tsallis–Jensen functional Tsallis and Jensen (2025) defined through a deformed Legendre structure, chosen so as to preserve a sensible canonical temperature on the Rindler patch. In the spherically symmetric sector the thermodynamic manifold is effectively one-dimensional, and the Clausius relation can be integrated to identify entropy and energy functionals under suitable regularity assumptions. The resulting equations of state describe an effective Newton coupling that becomes a function of area, together with emergent field equations that depart from the Einsteinian form while retaining a clear thermodynamic interpretation. Their analysis clarifies to what extent non-extensive entropies can be incorporated in an emergent-gravity setup without internal inconsistencies, but also makes manifest that many non-extensive models used in gravitational physics are, in effect, one-dimensional truncations where integrability constraints are relatively mild.
In this work, we follow a complementary and more conservative route. Rather than undertaking a full global analysis of horizon thermodynamics on a multi-parameter state space, we ask what can be inferred locally from a Jacobson-type construction once a non-extensive entropy slope is given. Concretely, we model the entropy of a horizon cross-section by the phenomenological power law $S(A)=\eta(A/4G)^{\delta}$, which is also known as Tsallis-$\delta$ entropy Tsallis and Cirto (2013), Barrow entropy Barrow (2020) or Barrow–Tsallis entropy Jizba et al. (2024). In its microcanonical form it also coincides with Tsallis–Jensen entropy Tsallis and Jensen (2025) which belong to the class of the so-called group entropies Tempesta (2016b, 2011, a); Jensen and Tempesta (2024). In the following, we will simply use the name power law entropy as it will better fit our purpose. With the power law entropy at hand, we will analyze how the corresponding entropy slope feeds into the effective Newton constant on local Rindler wedges.
@@ 2 Field equations from local horizon thermodynamics
In this context, non–additive entropies have been advocated as effective macroscopic descriptions for long-range interacting and gravitational systems, including black–hole thermodynamics and cosmology Caruso and Tsallis (2008); Tsallis and Cirto (2013); Saridakis et al. (2018). A particularly economical class is provided by Tsallis–Jensen group entropies, which preserve the first three Shannon–Khinchin axioms while incorporating a nontrivial composability structure in Tempesta’s sense Tempesta (2011, 2016a); Jensen and Tempesta (2024); Tsallis and Jensen (2025). Concretely, one considers functionals of the form
@@
$S_{R}^{(\alpha)}[p_{\rm eq}]\$$=$ $\ \frac{1}{1-\alpha}\,\ln\!\Bigg(\sum_{i=1}^{W(A)}p_{i}^{\alpha}\Bigg)\ =\ \frac{1}{1-\alpha}\,\ln\!\Big[W(A)\big(W(A)^{-1}\big)^{\alpha}\Big]$ (4)
$=$ $\ \frac{1}{1-\alpha}\,\ln\big(W(A)^{1-\alpha}\big)\ =\ \ln W(A)\,$,
@@
Evaluating the Tsallis–Jensen functional (2) on this microcanonical ensemble then yields
@@
where $I_{\rm m}[g,\psi]$ is the matter action. In particular, choosing $G_{\rm eff}=G$ and $\eta=1/(4G)$ reproduces the Wald entropy density for $f(R)$ gravity Wald (1993); Iyer and Wald (1994). This shows that the canonical Massieu framework is compatible with Jacobson’s non-equilibrium extension: an area-type entropy density leads directly to Einstein’s equations with $G_{\rm eff}=1/(4s_{0})$, while a curvature-dependent density $s(x)\propto f^{\prime}(R)$ supplemented by a suitable production term $d_{i}S$ reconstructs the field equations of $f(R)$ gravity in agreement with Eling et al. (2006); Brustein and Hadad (2009); Sotiriou and Faraoni (2010); De Felice and Tsujikawa (2010). In the following, we return to the area-type branch (11) and treat $s_{0}$ as the fundamental parameter controlling the local effective coupling via Eq. (49), namely $G_{\rm eff}=1/(4s_{0})$. The remaining task is to calibrate the resolution scale $A_{*}$, and hence $s_{0}$, in a non–arbitrary way, using only intrinsic data on the horizon cross section. In the next section we formulate this requirement as a Topological Calibration Principle, based on the Gauss–Bonnet relation between area, curvature and Euler characteristic, and use it to translate generalized entropy slopes into topology- and scale-dependent constraints on admissible non-extensive exponents.
@@ 3.1 Minimal calibration and reference area
Although our explicit formulas have focused on the phenomenologically simple power-law branch (11), the TCP logic relies only on the existence of a well-defined entropy slope $s_{0}=\partial S/\partial A$ evaluated at the calibrated area. For a more general functional $S(A,\chi)$, potentially with explicit topological dependence, the slope $s_{0}$ computed at $A_{*}=A_{0}(\chi)$ will, in general, depend on both $A_{*}$ and $\chi$, resulting in a topology-dependent $G_{\rm eff}$ of the same structural form as in Eq. (68). In this sense, the TCP imposes a self-consistency condition on how topology may enter generalized horizon entropies: viable non-extensive laws cannot be arbitrary functions of the area alone, but must encode topological information in a manner consistent with the Gauss–Bonnet calibration.

### v2

@@ 1 Introduction
A recent framework proposed by Lü, Di Gennaro and Ong  et al. (2025) takes an important step in this direction. In particular, working in a local-horizon framework in the spirit of Jacobson, they replace the Bekenstein–Hawking entropy by a stretched–exponential group–entropy functional Tempesta (2020); Tsallis and Jensen (2025) defined through a deformed Legendre structure, chosen so as to preserve a sensible canonical temperature on the Rindler patch. In the spherically symmetric sector the thermodynamic manifold is effectively one-dimensional, and the Clausius relation can be integrated to identify entropy and energy functionals under suitable regularity assumptions. The resulting equations of state describe an effective Newton coupling that becomes a function of area, together with emergent field equations that depart from the Einsteinian form while retaining a clear thermodynamic interpretation. Their analysis clarifies to what extent non-extensive entropies can be incorporated in an emergent-gravity setup without internal inconsistencies, but also makes manifest that many non-extensive models used in gravitational physics are, in effect, one-dimensional truncations where integrability constraints are relatively mild.
In this work, we follow a complementary and more conservative route. Rather than undertaking a full global analysis of horizon thermodynamics on a multi-parameter state space, we ask what can be inferred locally from a Jacobson-type construction once a non-extensive entropy slope is given. Concretely, we model the entropy of a horizon cross-section by the phenomenological power law $S(A)=\eta(A/4G)^{\delta}$, which is also known as Tsallis-$\delta$ entropy Tsallis and Cirto (2013), Barrow entropy Barrow (2020) or Barrow–Tsallis entropy Jizba et al. (2024). In its microcanonical form it also coincides with the stretched–exponential group entropy Tempesta (2020); Tsallis and Jensen (2025), which belongs to the class of the so-called group entropies Tempesta (2016b, 2011, a); Jensen and Tempesta (2024); Jensen et al. (2026). In the following, we will simply use the name power law entropy as it will better fit our purpose. With the power law entropy at hand, we will analyze how the corresponding entropy slope feeds into the effective Newton constant on local Rindler wedges.
@@ 2 Field equations from local horizon thermodynamics
In this context, non–additive entropies have been advocated as effective macroscopic descriptions for long-range interacting and gravitational systems, including black–hole thermodynamics and cosmology Caruso and Tsallis (2008); Tsallis and Cirto (2013); Saridakis et al. (2018). A particularly economical class is provided by Tempesta-type composable group entropies, which preserve the first three Shannon–Khinchin axioms while incorporating a nontrivial composability structure in Tempesta’s sense Tempesta (2020, 2011, 2016a); Jensen and Tempesta (2024); Tsallis and Jensen (2025). Concretely, one considers functionals of the form
@@
$S_{R}^{(\alpha)}[p_{\rm eq}]\$$=$ $\ \frac{1}{1-\alpha}\,\ln\!\Bigg(\sum_{i=1}^{W(A)}p_{i}^{\alpha}\Bigg)\ =\ \ln W(A)\,$, (4)
@@
Evaluating the group–entropy functional (2) on this microcanonical ensemble then yields
@@
where $I_{\rm m}[g,\psi]$ is the matter action. In particular, choosing $G_{\rm eff}=G$ and $\eta=1/(4G)$ reproduces the Wald entropy density for $f(R)$ gravity Wald (1993); Iyer and Wald (1994). This shows that the canonical Massieu framework is compatible with Jacobson’s non-equilibrium extension: an area-type entropy density leads directly to Einstein’s equations with $G_{\rm eff}=1/(4s_{0})$, while a curvature-dependent density $s(x)\propto f^{\prime}(R)$ supplemented by a suitable production term $d_{i}S$ reconstructs the field equations of $f(R)$ gravity in agreement with Eling et al. (2006); Brustein and Hadad (2009); Sotiriou and Faraoni (2010); De Felice and Tsujikawa (2010).
In the following, we return to the area-type branch (11) and treat $s_{0}$ as the fundamental parameter controlling the local effective coupling via Eq. (49), namely $G_{\rm eff}=1/(4s_{0})$. The remaining task is to calibrate the resolution scale $A_{*}$, and hence $s_{0}$, in a non–arbitrary way, using only intrinsic data on the horizon cross section. In the next section we formulate this requirement as a Topological Calibration Principle, based on the Gauss–Bonnet relation between area, curvature and Euler characteristic, and use it to translate generalized entropy slopes into topology- and scale-dependent constraints on admissible non-extensive exponents.